evaluate the integral
Evaluate the integral. [ 6x dy dx 0 - 1734 O 1020 0 - 204 O 102
Answer #1

landydami En dy da = 6x[9-6-83). = 6*117) = 102 x J 102x dx = 102 Judy 102a da 2102 [GI(40)] :=51(36-16) » Si(20) - 1020 as